OpenStudy (anonymous):

Simplify: 2x - 2y/ 3x + 3y times 6x + 6y/ x^2 - y^2

OpenStudy (btaylor):

\[\frac{2x-2y}{3x+3y} \times \frac{6x+6y}{x^2-y^2}\]\[\rightarrow \frac{(2)(6)(x-y)(x+y)}{(3)(x+y)(x+y)(x-y)}\]\[\rightarrow \frac{4}{x+y}\]

OpenStudy (anonymous):

thanks!